OpenStudy (anonymous):

an athlete completes one lap of a 400 m track in 54 s find her average velocity

OpenStudy (kropot72):

She runs 400 m in 54 seconds. How far does she run in 1 second?

OpenStudy (kropot72):

The distance that she runs in one second will give us the average velocity in meters per second.

OpenStudy (anonymous):

Average speed is 7.4m per second.

OpenStudy (kropot72):

Yes, you are correct. 400/54 = 7.4 m/s

OpenStudy (anonymous):

average velocity is zero

mathslover (mathslover):

@na2b4o7 is right! As the displacement becomes zero (the athlete returns to the same position from where she had started) . So : Average velocity = zero

OpenStudy (anonymous):

yes its 0

OpenStudy (anonymous):

average velocity=\[\frac{ Total displacement }{ Total time taken }\]

OpenStudy (kropot72):

It is true that the answer 7.4 m/s is the average speed and not the average velocity. However the statement in the question "....completes one lap......" does not give certainty that the runner returns to exactly the position of the starting point. Therefore a strictly correct answer is \[\bar{v}=\frac{d}{t}=\frac{vector\ displacement}{54}\ ms ^{-1}\]